OpenStudy (angiepangie0726):

Please help?? Fan and medal Solve 4x^2 − 24x + 31 = 0 by completing the square.

OpenStudy (luhvee1):

x =(24-√80)/8=3-1/2√ 5 = 1.882 x =(24+√80)/8=3+1/2√ 5 = 4.118

OpenStudy (angiepangie0726):

what do I do now?? @LuhVee1